OpenStudy (dls):

Find sum of last 4 digits of 3^3015.

OpenStudy (rational):

take mod 10^4

OpenStudy (dls):

can you explain how?

OpenStudy (rational):

maybe use binomial thm if you're not familiar wid congruences

OpenStudy (dls):

can you explain how? :p

OpenStudy (rational):

\[\large 3^{3015} = 3\cdot 3^{3014} = 3(3)^{2\times 1507} =3(9)^{ 1507} = 3(10-1)^{1507}\] expand

OpenStudy (dls):

I guess it will be 1507 only?

OpenStudy (dls):

I mean 1507*3

OpenStudy (rational):

im getting 8907

OpenStudy (dls):

\[C(1507,0)(10^{1507})({-1}^0) + C(1504,1)(10^{1506})(-1)^1+C(1504,2)(10^{1506})(-1)^2...\]

OpenStudy (zarkon):

go from the other direction

OpenStudy (dls):

\[ C(1507,0) (-1)^{1507}+ C(1507,1)({-1})^{1506}(10)+ C(1507,2)({-1})^{1505}(10)^2\]

OpenStudy (dls):

and so on

OpenStudy (rational):

so on until you get 10^4 term

OpenStudy (rational):

you can ignore all other terms because they dont affect last four digits

OpenStudy (dls):

I see..:O so it would be \[\Huge \sum_{k=1}^{k=4} C(1507,k)\] ?

OpenStudy (rational):

dont forget 3 dont forget alternate signs because of -1

OpenStudy (rational):

also the index must start from 0 and end at 3

OpenStudy (dls):

so..... \[\Huge \sum_{k=1}^{k=4} (-1)^kC(1507,k)\]

OpenStudy (dls):

oh yeah..0 to 3..the limits

OpenStudy (dls):

\[\Huge \sum_{k=0}^{k=3} (-1)^kC(1507,k)\]

OpenStudy (zarkon):

times 3

OpenStudy (dls):

yep!

OpenStudy (rational):

do we have \[(-1+10)^{1507} = \sum\limits_{k=0}^{1507} \binom{1507}{k}(-1)^{1507-k}10^k\] ?

OpenStudy (dls):

I got it..but I think the index should be from 0 to 4?

OpenStudy (rational):

0 to 3 will do

OpenStudy (dls):

10^0 ...10^1 ...10^2...10^3...10^4..till 10^4 there are 5 terms....

OpenStudy (rational):

there is no harm in taking more terms.. its just useless thats all

OpenStudy (zarkon):

there is harm

OpenStudy (rational):

10^4*(stuff) is divisible by 10^4 no harm

OpenStudy (dls):

do we have to take first 4 terms or till 10^4? earlier you said we gotta expand till 10^4...

OpenStudy (zarkon):

if you do mod 10^4 at each step you are ok...but if you don't then you will have problems

OpenStudy (zarkon):

there is subtraction in this problem

OpenStudy (rational):

expand till you get 10^4

OpenStudy (dls):

also the calculations are big in this expansion...without calculator its tedious..its an examination question where calculators usually aren't allowed so...

OpenStudy (rational):

10^4 mod 10^4 = 0 so the term that produces 10^4 as factor is not really required

OpenStudy (rational):

the moment you reach five digit number, you can ignore every thing except last four digits

OpenStudy (rational):

2309480293840 = 3840 in mod 10^4

OpenStudy (rational):

you can work first 3 binomial coefficients easily by hand i guess

OpenStudy (dls):

I see..that makes it crystal clear :D thanks a lot for your efforts :) @rational @Zarkon !

OpenStudy (rational):

you're not even allowed shop calculator for working factorials etc ?

OpenStudy (dls):

nope :|

OpenStudy (dls):

oh btw, C(1507,2) = 1134771 which is itself >10^4 ?

OpenStudy (rational):

then this problem should not be given in exam.. i see it is bit painful to work (1507 C 3) by hand because it is easy to make mistake while multiplying 1507*1506*1505 http://www.wolframalpha.com/input/?i=%283*%5Csum%5Climits_%7Bk%3D0%7D%5E3+%28%281507+choose+k%29%28-1%29%5E%281507-k%2910%5Ek%29%29+mod+10%5E4

OpenStudy (rational):

`C(1507,2) = 1134771 which is itself >10^4 ?` what good is this observation ?

OpenStudy (dls):

you said "he moment you reach five digit number, you can ignore every thing except last four digits"

OpenStudy (rational):

tell me this what are the last four digits of number 10^4*3498573498753498574398 + 0001 ?

OpenStudy (dls):

4399?

OpenStudy (rational):

sure ? look below : 10^4*3498573498753498574398 + 0001 34985734987534985743980000 + 0001

OpenStudy (dls):

oh..I see :P

OpenStudy (rational):

last four digits of 10^4*3498573498753498574398 + abcd is abcd

OpenStudy (dls):

yep.

OpenStudy (rational):

it is easy to convince ourself... clearly multiplying by 10^4 shifts all the digits to left by four decimal places producing four zeroes at the end.

OpenStudy (dls):

yep convinced..so..

OpenStudy (rational):

look at binomial expansion \[\begin{align}(-1+10)^{1507} &= \sum\limits_{k=0}^{1507} \binom{1507}{k}(-1)^{1507-k}10^k\\~\\&= \sum\limits_{k=0}^{3} \binom{1507}{k}(-1)^{1507-k}10^k+\color{blue}{ \sum\limits_{k=4}^{1507} \binom{1507}{k}(-1)^{1507-k}10^k}\end{align}\]

OpenStudy (rational):

you can factor out 10^4 from that blue part

OpenStudy (dls):

oh I see..so its not contributing..clear..:D